It was known that the actual length of the beam was 6.3 in. The class was asked to measure the beam using different types of measuring tapes. Demetrius's tape was accurate to one-eighth of an inch, so he measured the board to be 6.25 in. What was the percent error of his measurement to the nearest tenth of a percent?

Answer by Theo(13342)

actual length of beam = 6.3
measured length of beam = 6.25
error was 6.3 - 6.25 = .05 inches.
eror ratio = .05/6.3 = .0079365079 * 100 = .79365079 percent.
round to nearest tenth of a percent = .8 percent.
percent error = .8 is your solution.